# Umlaut information

### v1

@@ III.2 Closed form and optimal marginal
Let us call in general $\ddot{P}_{Y}^{(\alpha)}$ the $\alpha$-umlaut marginal of an arbitrary distribution $P_{XY}$, as in (42). If we consider any product distribution $P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}}$, it is easy to verify that its $\alpha$-umlaut reduced state is the tensor product of $\ddot{P}_{Y}$ and $\ddot{P}^{\prime}_{Y^{\prime}}$. Since the Rényi quantum relative entropies are additive, this concludes the proof of the additivity of $U_{\alpha}$. ∎
III.3 Operational interpretation in composite hypothesis testing
@@
[…]
@@
$D^{\varepsilon}_{H}(Q^{\times n}\|P^{\times n})\coloneqq-\log\min\left\{\beta(\mathcal{A}_{n})\,:\,\alpha(\mathcal{A}_{n})\leq\varepsilon\right\}.$ (45)
The type II-error exponent will asymptotically decay according to the error exponent
$\text{Stein}(Q\|P)\coloneqq\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(Q^{\times n}\|P^{\times n})\,,$ (46)
which we refer to as Stein’s exponent. Stein’s lemma [15, 16] states that
$\text{Stein}(Q\|P)=D(Q\|P)\,,$ (47)
@@
Here, we are interested in the generalisation of the above setting where the null hypothesis is composite, i.e. non-i.i.d. Namely, for all $n\in\mathbb{N}^{+}$, let $\mathcal{F}_{n}\subseteq\mathcal{P}(\mathcal{X}^{n})$ be a given family of distributions; we collect the sets $\mathcal{F}_{n}$ in the sequence $\mathcal{F}=(\mathcal{F}_{n})_{n\in\mathbb{N}^{+}}$. The new task is testing whether $X^{n}$ is distributed according to $Q_{n}$, for some $Q_{n}\in\mathcal{F}_{n}$, or according to the i.i.d. distribution $P^{\times n}$. The error probabilities are
[…]
Minimising the type II-error with the usual constraint on the type I-error yields
$\mathrm{Stein}(\mathcal{F}\|P)\coloneqq\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n}),$ (50)
@@
$D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n})\coloneqq-\log\min\left\{\beta(\mathcal{A}_{n})\,:\,\alpha(\mathcal{A}_{n})\leq\varepsilon\right\},$ (51)
@@
$D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n})=\min_{Q_{n}\in\mathcal{F}_{n}}D^{\varepsilon}_{H}(Q_{n}\|P^{\times n})\,,$ (52)
@@
The operational interpretation of the umlaut information emerges in the composite hypothesis testing problem where the underlying alphabet is bipartite, i.e. of the form $\mathcal{X}\times\mathcal{Y}$, the alternative hypothesis is i.i.d. with single-copy distribution $P\mapsto P_{XY}$, and the null hypothesis is composite and of the form $Q_{n}\mapsto P_{X}^{\times n}Q_{Y^{n}}$ [18, 19]. Here, $P_{X}$ is the marginal of $P_{XY}$ to the $X$ variable, and $Q_{Y^{n}}$ denotes an arbitrary probability distribution on $\mathcal{Y}^{n}$.
@@
$\mathcal{F}^{P_{X}}_{n}$$\coloneqq\left\{P_{X}^{\times n}Q_{Y^{n}}:\ Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y^{n}})\right\}.$ (53)
@@
$D^{\varepsilon}_{H}\big{(}\mathcal{F}_{n}^{P_{X}}\,\big{\|}\,P_{XY}^{\times n}\big{)}=\min_{Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y^{n}})}D^{\varepsilon}_{H}\big{(}P_{X}^{\times n}Q_{Y^{n}}\,\big{\|}\,P_{XY}^{\times n}\big{)}\,.$ (54)
The following theorem states that the corresponding Stein exponent coincides with the umlaut information between $X$ and $Y$.
@@
Given a joint probability distribution $P_{XY}\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$, let $P_{X}$ be the marginal on $\mathcal{X}$. Then it holds that $U(X;Y)=\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}\,.$ (55)
@@
$P^{\times n}_{XY}(x_{1},\dots,x_{n},y_{1},\dots,y_{n})=P_{XY}(x_{1},y_{1})\cdots P_{XY}(x_{n},y_{n})\,.$ (56)
@@
$D^{\varepsilon}_{H}(p\|q)\geq D_{\alpha}(p\|q)+\frac{\alpha}{1-\alpha}\log\frac{1}{\varepsilon}.$ (57)
@@
$\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ =$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}\min_{Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y}^{n})}D^{\varepsilon}_{H}(P_{X}^{\times n}Q_{Y^{n}}\|P_{XY}^{\times n})$ (58)
@@
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}U_{\alpha}(X;Y),$
@@
$\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ $\geq\limsup_{\alpha\to 1^{-}}U_{\alpha}(X;Y)$ (59)
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}U(X;Y),$
@@
$Q_{Y^{n}}(y_{1},\dots,y_{n})=Q_{Y}^{\times n}(y_{1},\ldots,y_{n})=Q_{Y}(y_{1})\cdots Q_{Y}(y_{1}),$ (60)
where $Q_{Y}\in\mathcal{P}(\mathcal{Y})$ is an arbitrary fixed distribution. Then, continuing from the first line of (58) we have
$\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ $\leq\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(P_{X}^{n}Q_{Y}^{n}\|P_{XY}^{n})$ (61)
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iv)}}}}{{=}}D(P_{X}Q_{Y}\|P_{XY})$
where (iv) follows from the Stein’s lemma [15, 16]. Minimising over $Q_{Y}\in\mathcal{P}(\mathcal{Y})$ yields
$\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}\leq\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})=U(X;Y)\,,$ (62)
@@
III.4 Example: Joint Gaussian distributions
The umlaut information (Definition 2) also applies to continuous variables and an instructive example is follows.

### v2

@@ III.3 Rényi divergence variant
Let us call in general $\ddot{P}_{Y}^{(\alpha)}$ the $\alpha$-umlaut marginal of an arbitrary distribution $P_{XY}$, as in (45). If we consider any product distribution $P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}}$, it is easy to verify that its $\alpha$-umlaut reduced state is the tensor product of $\ddot{P}_{Y}$ and $\ddot{P}^{\prime}_{Y^{\prime}}$. Since the Rényi quantum relative entropies are additive, this concludes the proof of the additivity of $U_{\alpha}$.
We now consider the case of $\alpha>1$. Let us assume $U_{\alpha}(X;Y)<\infty$.
$\hskip 0.0ptU_{\alpha}(X;Y)$ $=\min_{Q_{Y}}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)P_{XY}^{1-\alpha}(x,y)$ (47)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y}^{\ast})}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathrm{supp}(P_{X})\\ y\in\mathcal{Y}^{\ast}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)P_{XY}^{1-\alpha}(x,y)$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y}^{\ast})}\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}^{\ast}}Q_{Y}^{\alpha}(y)\left(\frac{1}{Z_{\alpha}[P_{XY}]}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}\right)^{1-\alpha}-\log Z_{\alpha}[P_{XY}]$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y}^{\ast})}D_{\alpha}(Q_{Y}\|\ddot{P}_{Y}^{(\alpha)})-\log Z_{\alpha}[P_{XY}]$
where, starting from (i), we denote $\mathcal{Y}^{\ast}\coloneqq\{y\in\mathcal{Y}:P(x,y)>0\ \ \forall x\in\mathrm{supp}(P_{X})\}$; we can indeed restrict the optimization to probability distributions supported on $\mathcal{Y}^{\ast}$ since, if $Q_{Y}(y)>0$ for $y\in\mathcal{Y}\setminus\mathcal{Y}^{\ast}$, then we would have
$\hskip 0.0pt\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)P_{XY}^{1-\alpha}(x,y)=+\infty.$ (48)
Clearly, $\mathcal{Y}^{\ast}=\emptyset$ if and only if $U_{\alpha}(X;Y)=+\infty$. In (ii) we have introduced
$\hskip 0.0ptZ_{\alpha}[P_{XY}]\coloneqq\sum_{y\in\mathcal{Y}^{\ast}}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}$ (49)
and in (iii) we have defined
$\hskip 0.0pt\ddot{P}_{Y}^{(\alpha)}(y)\coloneqq\frac{1}{Z_{\alpha}[P_{XY}]}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}\qquad y\in\mathcal{Y}^{\ast},$ (50)
which is a probability distribution on $\mathcal{Y}^{\ast}$ and which can naturally be extended to a probability distribution on $\mathcal{Y}$ by setting $\ddot{P}_{Y}^{(\alpha)}(y)=0$ for all $y\in\mathcal{Y}\setminus\mathcal{Y}^{\ast}$. Using the conventions $0^{-1}=+\infty$ $\infty^{-1}=0$, since $\alpha>1$ we could formally rewrite
$\hskip 0.0ptZ_{\alpha}[P_{XY}]=\sum_{y\in\mathcal{Y}}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}},$ (51)
extending the sum to $\mathcal{Y}$, and similarly
$\hskip 0.0pt\ddot{P}_{Y}^{(\alpha)}(y)\coloneqq\frac{1}{Z_{\alpha}[P_{XY}]}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}},\qquad y\in\mathcal{Y},$ (52)
Looking at the last line of (47), we notice that
$\hskip 0.0ptU_{\alpha}(X;Y)=-\log Z_{\alpha}[P_{XY}]$ (53)
and that the minimisation problem is saturated for $Q_{Y}=\ddot{P}_{Y}^{(\alpha)}$, which concludes the proof of (40) and (41). If we consider $P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}}$ and we construct
$\hskip 0.0pt(\mathcal{Y}\times\mathcal{Y}^{\prime})^{\ast}=\{(y,y^{\prime})\in\mathcal{Y}\times\mathcal{Y}^{\prime}:P_{XY}(x,y)P^{\prime}_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})>0\ \ \forall(x,x^{\prime})\in\mathrm{supp}(P_{X}P^{\prime}_{X^{\prime}})\},$ (54)
then $(\mathcal{Y}\times\mathcal{Y}^{\prime})^{\ast}=\mathcal{Y}^{\ast}\times(\mathcal{Y}^{\prime})^{\ast}$. Indeed,
[…]
So, the optimiser in $U_{\alpha}(XX^{\prime};YY^{\prime})$ is proportional to
$\hskip 0.0pt\mathbbm{1}_{(\mathcal{Y}\times\mathcal{Y}^{\prime})^{\ast}}(y,y^{\prime})$ $\left(\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ x^{\prime}\in\mathcal{X}^{\prime}\end{subarray}}P_{X}^{\alpha}(x)P_{X^{\prime}}^{\alpha}(x^{\prime})P_{XY}^{1-\alpha}(x,y){P^{\prime}}_{X^{\prime}Y^{\prime}}^{1-\alpha}(x^{\prime},y^{\prime})\right)^{\frac{1}{1-\alpha}}$ (56)
$=\mathbbm{1}_{\mathcal{Y}^{\ast}}(y)\mathbbm{1}_{(\mathcal{Y}^{\prime})^{\ast}}(y^{\prime})\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}\left(\sum_{x^{\prime}\in\mathcal{X}^{\prime}}{P^{\prime}}_{X^{\prime}}^{\alpha}(x^{\prime})P_{XY}^{1-\alpha}(x^{\prime},y^{\prime})\right)^{\frac{1}{1-\alpha}},$
whence
$\hskip 0.0ptU_{\alpha}(XX^{\prime};YY^{\prime})$ $=D_{\alpha}(P_{X}\ddot{P}_{Y}^{(\alpha)}P^{\prime}_{X^{\prime}}\ddot{P^{\prime}}_{Y^{\prime}}^{(\alpha)}\|P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}})$ (57)
$=D_{\alpha}(P_{X}\ddot{P}_{Y}^{(\alpha)}\|P_{XY})+D_{\alpha}(P^{\prime}_{X^{\prime}}\ddot{P^{\prime}}_{Y^{\prime}}^{(\alpha)}\|P^{\prime}_{X^{\prime}Y^{\prime}})=U_{\alpha}(X;Y)+U_{\alpha}(X^{\prime};Y^{\prime}).$
This concludes the proof of the additivity.
Remark 11.
For $\alpha\in(0,1)\cup(1,\infty)$, we can rewrite, similarly to (32),
$\hskip 0.0ptU_{\alpha}(X;Y)$ $=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{X|Y}^{1-\alpha}(x|y)\right)^{\frac{1}{1-\alpha}}$ (58)
$=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\exp\left(-D_{\alpha}(P_{X}\|P_{X|Y=y})\right)$
$=-\log\mathbb{E}_{Z\sim P_{Y}}\left[\exp\left(-D_{\alpha}(P_{X}\|P_{X|Y=Z})\right)\right].$
This form suggests another proof of Lemma 9, where we showed that the Rényi $\alpha$-umlaut information converges to the umlaut information in the limit $\alpha\to 1$.
Alternative proof of Lemma 9.
By the closed-form expression in (58), we can compute
$\hskip 0.0pt\lim_{\alpha\to 1^{\pm}}U_{\alpha}(X;Y)$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\exp\left(-\lim_{\alpha\to 1^{\pm}}D_{\alpha}(P_{X}\|P_{X|Y=y})\right)$ (59)
$=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\exp\left(-D(P_{X}\|P_{X|Y=y})\right)$
$=U(X;Y).$
where in (i) we have used that the sum is over a finite set. ∎
III.4 Operational interpretation in composite hypothesis testing
@@
$\alpha(\mathcal{A}_{n})$ $\coloneqq\mathbb{P}_{X^{n}\sim Q^{\times n}}\left(\mathcal{A}_{n}(X_{1},\dots,X_{n})=H_{1}\right)$ $\text{type-I error},$ (60)
$\beta(\mathcal{A}_{n})$ $\coloneqq\mathbb{P}_{X^{n}\sim P^{\times n}}\left(\mathcal{A}_{n}(X_{1},\dots,X_{n})=H_{0}\right)$ $\text{type-II error}.$ (61)
@@
$\hskip 0.0ptD^{\varepsilon}_{H}(Q^{\times n}\|P^{\times n})\coloneqq-\log\min\left\{\beta(\mathcal{A}_{n})\,:\,\alpha(\mathcal{A}_{n})\leq\varepsilon\right\}.$ (62)
The type II error exponent will asymptotically decay according to the error exponent
$\hskip 0.0pt\text{Stein}(Q\|P)\coloneqq\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(Q^{\times n}\|P^{\times n})\,$, (63)
which we refer to as Stein exponent. The Chernoff–Stein lemma [16, 17] states that
$\hskip 0.0pt\text{Stein}(Q\|P)=D(Q\|P)\,$, (64)
@@
Here, we are interested in the generalisation of the above setting where the null hypothesis is composite, i.e. is given by a set of non-i.i.d distributions. Namely, for all $n\in\mathbb{N}^{+}$, let $\mathcal{F}_{n}\subseteq\mathcal{P}(\mathcal{X}^{n})$ be a given family of distributions; we collect the sets $\mathcal{F}_{n}$ in the sequence $\mathcal{F}=(\mathcal{F}_{n})_{n\in\mathbb{N}^{+}}$. The new task is testing whether $X^{n}$ is distributed according to $Q_{n}$, for some $Q_{n}\in\mathcal{F}_{n}$, or according to the i.i.d. distribution $P^{\times n}$. The error probabilities are
$\alpha(\mathcal{A}_{n})$ $\coloneqq\sup_{Q_{n}\in\mathcal{F}_{n}}\mathbb{P}_{X^{n}\sim Q_{n}}\left(\mathcal{A}_{n}(X_{1},\dots,X_{n})=H_{1}\right)$ $\text{type-I error},$ (65)
$\beta(\mathcal{A}_{n})$ $\coloneqq\mathbb{P}_{X^{n}\sim P^{\times n}}\left(\mathcal{A}_{n}(X_{1},\dots,X_{n})=H_{0}\right)$ $\text{type-II error}.$ (66)
Minimising the type II error with the usual constraint on the type I error yields a decay rate for the former equal to
$\hskip 0.0pt\mathrm{Stein}(\mathcal{F}\|P)\coloneqq\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n}),$ (67)
@@
$\hskip 0.0ptD^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n})\coloneqq-\log\min\left\{\beta(\mathcal{A}_{n})\,:\,\alpha(\mathcal{A}_{n})\leq\varepsilon\right\},$ (68)
@@
$\hskip 0.0ptD^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n})=\min_{Q_{n}\in\mathcal{F}_{n}}D^{\varepsilon}_{H}(Q_{n}\|P^{\times n})\,$, (69)
@@
We can also consider the strong converse Stein exponent, given by a modified version of (67) in which we allow the type I error to be arbitrarily close to $1$, instead of arbitrarily small, and we replace the limit inferior in $n$ with a limit superior:
$\hskip 0.0pt\mathrm{Stein}^{\dagger}(\mathcal{F}\|P)\coloneqq\lim_{\varepsilon\to 1^{-}}\limsup_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n}).$ (70)
Clearly, in general $\mathrm{Stein}^{\dagger}(\mathcal{F}\|P)\geq\mathrm{Stein}(\mathcal{F}\|P)$. In many interesting cases, however, equality holds; equivalently, $\lim_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(\mathcal{F}_{n}\|P^{\times n})$ exists for all $\varepsilon\in(0,1)$, and its value is independent of $\varepsilon$. This holds, for example, when both hypotheses are simple and i.i.d., in which case we have indeed [16, 17]
$\hskip 0.0ptD(Q\|P)=\mathrm{Stein}(Q\|P)=\mathrm{Stein}^{\dagger}(Q\|P)=\lim_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(Q^{\times n}\|P^{\times n})\qquad\forall\ \varepsilon\in(0,1)\,.$ (71)
The operational interpretation of the umlaut information emerges in the composite hypothesis testing problem where the underlying alphabet is bipartite, i.e. of the form $\mathcal{X}\times\mathcal{Y}$, the alternative hypothesis is i.i.d. with single-copy distribution $P\mapsto P_{XY}$, and the null hypothesis is composite and of the form $Q_{n}\mapsto P_{X}^{\times n}Q_{Y^{n}}$. Here, $P_{X}$ is the marginal of $P_{XY}$ to the $X$ variable, and $Q_{Y^{n}}$ denotes an arbitrary probability distribution on $\mathcal{Y}^{n}$. With the role of the two hypotheses interchanged, this problem previously appeared in [19, 20, 21].
@@
$\hskip 0.0pt\mathcal{F}^{P_{X}}_{n}$ $\coloneqq\left\{P_{X}^{\times n}Q_{Y^{n}}:\ Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y}^{n})\right\}.$ (72)
@@
$\hskip 0.0ptD^{\varepsilon}_{H}\big{(}\mathcal{F}_{n}^{P_{X}}\,\big{\|}\,P_{XY}^{\times n}\big{)}=\min_{Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y^{n}})}D^{\varepsilon}_{H}\big{(}P_{X}^{\times n}Q_{Y^{n}}\,\big{\|}\,P_{XY}^{\times n}\big{)}\,.$ (73)
The following theorem states that the corresponding Stein exponent coincides with the umlaut information between $X$ and $Y$, and that this equality holds also in the strong converse regime.
@@
Given a joint probability distribution $P_{XY}\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$, let $P_{X}$ be the marginal on $\mathcal{X}$. Then it holds that $\hskip 0.0ptU(X;Y)=\lim_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}\qquad\forall\ \varepsilon\in(0,1)\,;$ (74) equivalently, $\hskip 0.0ptU(X;Y)=\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}=\mathrm{Stein}^{\dagger}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}\,.$ (75)
@@
$\hskip 0.0ptP^{\times n}_{XY}(x_{1},\dots,x_{n},y_{1},\dots,y_{n})=P_{XY}(x_{1},y_{1})\cdots P_{XY}(x_{n},y_{n})\,.$ (76)
@@
$\hskip 0.0ptD^{\varepsilon}_{H}(p\|q)\geq D_{\alpha}(p\|q)+\frac{\alpha}{1-\alpha}\log\frac{1}{\varepsilon}.$ (77)
@@
$\hskip 0.0pt\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}\min_{Q_{Y^{n}}\in\mathcal{P}(\mathcal{Y}^{n})}D^{\varepsilon}_{H}(P_{X}^{\times n}Q_{Y^{n}}\|P_{XY}^{\times n})$ (78)
@@
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}U_{\alpha}(X;Y),$
@@
$\hskip 0.0pt\mathrm{Stein}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ $\geq\lim_{\alpha\to 1^{-}}U_{\alpha}(X;Y)$ (79)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}U(X;Y),$
@@
$\hskip 0.0ptQ_{Y^{n}}(y_{1},\dots,y_{n})=Q_{Y}^{\times n}(y_{1},\ldots,y_{n})=Q_{Y}(y_{1})\cdots Q_{Y}(y_{1}),$ (80)
where $Q_{Y}\in\mathcal{P}(\mathcal{Y})$ is an arbitrary fixed distribution. Then,
$\hskip 0.0pt\mathrm{Stein}^{\dagger}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}$ $\leq\lim_{\varepsilon\to 1^{-}}\limsup_{n\to\infty}\frac{1}{n}D^{\varepsilon}_{H}(P_{X}^{n}Q_{Y}^{n}\|P_{XY}^{n})$ (81)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iv)}}}}{{=}}D(P_{X}Q_{Y}\|P_{XY})$
where (iv) follows from the strong converse of the Chernoff–Stein lemma [16, 17], here reported as (71). Minimising over $Q_{Y}\in\mathcal{P}(\mathcal{Y})$ yields
$\hskip 0.0pt\mathrm{Stein}^{\dagger}\big{(}\mathcal{F}^{P_{X}}\,\big{\|}\,P_{XY}\big{)}\leq\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})=U(X;Y)\,$, (82)
@@
III.5 Example: Joint Gaussian distributions
The definition of the umlaut information (Definition 2) also applies to continuous variables (see Appendix B), and an instructive example is as follows.